OpenStudy (anonymous):

Trying to solve 2/16×3/4 to the lowest form

OpenStudy (anonymous):

you have to get a least common denominator

OpenStudy (anonymous):

simplify the fractions and they will be 1/8*3/4

OpenStudy (anonymous):

you don't need common denomitator, thats for addition and subtraction

OpenStudy (anonymous):

you just multipli numeraor with numerator and denomitator with denominator

OpenStudy (anonymous):

multiply*

OpenStudy (anonymous):

now, \[\frac{ 1 }{ 8 } \times \frac{ 3 }{ 4 }\] if you have that then multiply nmerator times numerator and denominator times denominator, which will be \[\frac{ 3 }{ 32 }\]